How many five-card hands consisting of three cards of one face value and two cards of another face value can be dealt from a deck of 48 playing cards?
Answer by stanbon(75887)

How many five-card hands consisting of three cards of one face value and two cards of another face value can be dealt from a deck of 48 playing cards?
------------
----------------------------
# of ways to select 3 cards of one face value: 13*4C3 = 13*4 = 52
# of ways to select 2 cards of another face value: 12*4C2 = 72
---
# of ways to "succeed":: 52*72
----
# of random 5-card hands:: 52C5
-------
Ans: P(condition) = (52*72)/52C5 = 0.00144
------------
Cheers,
Stan H.
------------